ALCOMA ’05 Thurnau (Germany) April 03-10, 2005

Registration Form

Please complete one form for each delegate. Return each completed form to:

ALCOMA 05

Lehrstuhl II f. Mathematik

Universität Bayreuth

D-95440 Bayreuth

Germany Fax: 0049-(0)921-55-3385

1. Delegate Information:

Title: First name:

Last name:

Postal Address:

Email address:

2. Participation Fee:

The participation fee is EURO 110, payable at the conference. There are no fees for accompanying persons.

3. Contributed Talks:

Do you intend to present a contributed talk: □ yes □ no

Title of the talk:

Please send your abstract, preferably by email in L^{A}TEX-Format, in the style specified in the accompanying instructions by January 09, 2005.

4. Hotel Accomodation (will be provided by the local tourist office):

I need accomodation for persons.

Date of arrival:  date of departure:  .

Date: Signature: